# Compliance-to-Code: Enhancing Financial Compliance Checking via Code Generation

### v2

@@ 6.1 Task 1 Performance: Advancing Automated Structuring of Complex Regulatory Text
Table 1: Task 1 Performance (Precision, Recall, F1 %) on the Test Set. CU Fields denotes average scores for Subject, Condition, Constraint. CU Identification concerns overall span detection. Best fine-tuned model scores are in bold. SFT indicates Supervised Fine-Tuning.
@@
DeepSeek-V3 71.05 68.01 69.50 49.11 49.11 49.11
Claude-3.7 $65.28$ $64.49$ $64.89$ $41.07$ $41.07$ $41.07$
@@
GLM-4-9B-0414 (SFT) 72.10 71.27 71.68 37.50 37.50 37.50
Qwen3-8B (SFT) $63.45$ $63.03$ $63.24$ $31.25$ $31.25$ $31.25$
DeepSeek-R1-Distill-Qwen-7B (SFT) $54.97$ $53.28$ $54.11$ $23.21$ $23.21$ $23.21$
Our evaluation demonstrates that domain-adapted supervised fine-tuning is indispensable for structuring complex regulatory text into actionable CUs and inter-unit relations. The fine-tuned GLM-4-9B-0414 achieves 71.7% F1 in CU field extraction, outperforming strong few-shot baselines such as GPT-4.1 by nearly 8 absolute F1 points.
Error analysis localizes major bottlenecks in nuanced relation disambiguation (e.g., should include vs refer to), and multi-hop dependencies. Despite these, the structuring phase delivers a reliable foundation for downstream code synthesis, preserving the majority of regulatory semantics and enabling scalable automation for dense financial texts.
@@
Direct Prompting Baselines (Task 3: Regulation-to-Code, Model: Qwen3-8B)
@@
Direct Prompting Baselines (Task 2: Structure-to-Code, Model: Qwen3-8B, Gold Struct)
@@
Direct Prompting Baselines (Task 2: Structure-Reasoning-to-Code, Model: Qwen3-8B, Gold Struct)
@@ 6.2 Tasks 2&3: Structured Regulations Are Transformative for Compliance Code Generation
Table 3 demonstrates that structured input and explicit reasoning steps substantially improve code generation performance. Moving from direct prompting on raw regulations (Task 3, Pass@1 $\sim$52.9%) to providing human-annotated structures (Task 2, Pass@1 $\sim$61.0%) yields a clear performance boost. Incorporating explicit reasoning (Task 2 Structure-Reasoning-to-Code) further elevates Pass@1 to 78.1%. This progression highlights the importance of decomposing complex compliance requirements into structured representations and guiding the model through intermediate reasoning steps.
@@ 7 Conclusion
We have introduced the Compliance-to-Code Dataset and a novel structured synthesis pipeline, embodied in our FinCheck system, for executable compliance checking in the complex domain of financial regulation. Our approach, which formalizes regulatory logic as annotated Compliance Units and translates these into verifiable Python code, demonstrates significant gains in functional accuracy (achieving up to 78.1% Pass@1 with gold structures and reasoning steps, interpretability through traceable commentary, and inherent auditability compared to end-to-end LLM prompting. This research not only provides a valuable new resource and strong benchmark results but also opens new avenues for developing safer, more scalable, and reliable RegTech solutions.
@@
Societal Impact: Our approach can streamline compliance and enhance transparency, but should not replace expert judgment in critical situations. We have carefully reviewed and redacted data for release, and recommend that our tools be used with responsible human oversight. For more discussion, please refer to Ethics Statements in Appendix D.
Appendix A License
The Compliance-to-Code dataset will be publicly available and use the CC-BY-NC-4.0 license.
@@ Appendix B Accessibility
Appendix C Data Usage
@@
Appendix D Ethics Statements
The research conducted in this paper adheres in every respect to the NeurIPS Code of Ethics.
@@ Appendix D Ethics Statements
Appendix E Acknowledgements
We thank Siyuan Li for leading the dataset construction and experiments, ensuring data quality and experimental accuracy. Jian Chen served as the project lead and contributed to the overall research framework, dataset and experimental design, manuscript preparation, and initial project coordination. Rui Yao contributed to both dataset construction and experimental implementation. Weihua Qiu, Simin Zhang, Chucheng Dong, and Zhiyao Li assisted in dataset creation. Xuming Hu and Peilin Zhou contributed to the revision and optimization of the manuscript. We thank Qipeng Xie and Zixuan Yuan for providing computational resources, with Zixuan Yuan also offering academic guidance. We are grateful to Dr. Yutao Huang (Chinese Academy of Social Sciences) for policy consulting, and to Yuting Qiu for legal advice.
This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0630) and NSFC No. 6240074265.
@@ Appendix H Prompt of Instruction-tuning
Because the length of prompt is long, we only display the structure of the prompt in Figure 4 For more detail, please refer to the dataset.

### v3

@@ 6.1 Task 1 Performance: Advancing Automated Structuring of Complex Regulatory Text
Table 1: Task 1 Performance (Precision, Recall, F1 %) on the Test Set. CU Fields denotes average scores for Subject, Condition, Constraint. CU Identification concerns overall span detection. Best model scores are in bold. All Fine-tuned Models are performed using Supervised Fine-Tuning (SFT).
@@
DeepSeek-V3-0324 $71.05$ $68.01$ $69.50$ $49.11$ $49.11$ $49.11$
Claude-3.7-Sonnet $65.28$ $64.49$ $64.89$ $41.07$ $41.07$ $41.07$
Gemini-2.5-pro $71.33$ $70.59$ $70.96$ $48.21$ $48.21$ $48.21$
DeepSeek-R1-0528 72.01 70.97 71.49 50.00 50.00 50.00
Fin-R1 $58.08$ $53.92$ $55.92$ $13.39$ $13.39$ $13.39$
@@
GLM-4-9B-0414 72.10 71.27 71.68 37.50 37.50 37.50
Qwen3-8B $63.45$ $63.03$ $63.24$ $31.25$ $31.25$ $31.25$
DeepSeek-R1-Distill-Qwen-7B $54.97$ $53.28$ $54.11$ $23.21$ $23.21$ $23.21$
Our results indicate that domain-specific supervised fine-tuning (SFT) is an effective strategy. This is demonstrated by the GLM-4-9B-0414 model. As a 9B parameter model, its performance in CU field extraction after SFT (71.68% F1) surpasses that of powerful proprietary models in a few-shot setting, including DeepSeek-R1-0528 (71.49% F1) and Gemini-2.5-pro (70.96% F1). This outcome suggests that a smaller, well-tuned model can outperform larger, general-purpose counterparts on specialized tasks, representing an efficient path to achieving a high level of performance.
@@
Fin-R1 $37.92$ $62.74$ $68.03$ $50.31$ $77.64$ $75.11$ $12.94$ $36.52$ $54.50$ $5.82$ $21.77$ $49.21$
DeepSeek-R1-0528 73.67 - 73.34 81.04 - 80.20 70.59 - 63.76 47.27 - 52.96
Error analysis pinpoints a primary bottleneck in the propagation of contextual information, particularly within complex legal obligations that span multiple sentences. Models often struggle to maintain and transmit contextual cues across these long-form dependencies, leading to omissions or misinterpretations. Despite these remaining challenges, the structuring phase delivers a reliable foundation for downstream code synthesis, preserving the majority of regulatory semantics and enabling scalable automation for dense financial texts.
@@
Prompting Baselines (Task 3: Regulation-to-Code, Model: Qwen3-8B)
@@
Prompting Baselines (Task 2: Structure-to-Code, Model: Qwen3-8B, Gold Struct)
@@
Composed Pipeline (Regulation-to-Structure-to-Code, Model: Qwen3-8B, Predicted Struct)
CoT-LLM (one-shot) 54.58 69.39 69.12 70.55 87.64 77.24 43.53 66.87 55.62 0.18 0.91 46.33
Reasoning-Augmented Setting (Structure-Reasoning-to-Code, Model: Qwen3-8B, Gold Struct)
@@ 6.2 Tasks 2&3: Structured Regulations Are Transformative for Compliance Code Generation
Table 3 demonstrates that structured input and explicit reasoning steps substantially improve code generation performance. Moving from direct prompting on raw regulations (Task 3, Pass@1 $\sim$52.9%) to providing human-annotated structures (Task 2, Pass@1 $\sim$61.0%) yields a clear performance boost. Incorporating explicit reasoning (Structure-Reasoning-to-Code) further elevates Pass@1 to 78.1%. This progression highlights the importance of decomposing complex compliance requirements into structured representations and guiding the model through intermediate reasoning steps. We further investigated in Regulation-to-Structure-to-Code. This pipeline shows an improvement over the Regulation-to-Code baseline, while the performance on "Difficult" tasks does not improve significantly. This underscores a critical challenge: error propagation. Upstream structuring errors propagate to the code generation stage, exposing model limitations in complex, multi-hop legal reasoning.
@@
6.3 Trade-off Between Compliance Reasoning and Stylistic Alignment
Epoch-wise Divergence of SFT
Results in Table 4 illustrate a clear divergence between stylistic alignment and compliance reasoning during extended LoRA-based SFT of Qwen3-8B on structure-to-code tasks. Increasing epochs consistently enhance CodeBLEU, reflecting improved mimicry of human coding patterns, but fail to yield corresponding improvements in compliance reasoning as evidenced by declining Pass@1. The simultaneous moderate gain in Pass@5 suggests higher output diversity yet degraded confidence calibration. We conjecture that this phenomenon stems from catastrophic forgetting intertwined with over-adaptation: prolonged training incrementally biases the model toward superficial pattern memorization, consequently compromising deeper logical inference capability. These observations highlight an inherent tension in fine-tuning strategies, suggesting a critical need for regularization techniques or strategic early stopping to preserve compliance reasoning amid stylistic improvements.
Table 4: Performance Evaluation of Qwen3-8B across multiple SFT epochs on Code Generation Tasks. 20% of the original training set was divided into validation set to monitor performance. Best results are highlighted in bold.
Model Overall (N=307) Simple (N=215) Medium (N=34) Difficult (N=58)
Fine-tuned Model
Qwen3-8B-1-epoch 56.27 80.74 $69.93$ 66.18 88.37 $77.33$ 41.18 69.14 $56.98$ 27.55 $58.64$ $49.55$
Qwen3-8B-3-epoch $52.18$ $76.39$ $69.99$ $65.05$ $84.99$ $77.90$ $28.53$ $65.56$ $57.23$ $17.45$ $50.10$ $47.54$
Qwen3-8B-5-epoch $49.38$ $75.99$ $70.52$ $62.37$ $85.72$ $78.04$ $19.41$ $54.97$ $57.48$ $18.09$ $51.63$ $49.74$
Qwen3-8B-7-epoch $50.35$ $78.38$ $71.10$ $62.63$ $87.91$ 78.53 $21.18$ $54.20$ $58.74$ $21.27$ $56.77$ $50.20$
Qwen3-8B-9-epoch $51.10$ $80.39$ 71.30 $63.48$ $87.84$ $78.07$ $24.71$ $64.74$ 61.11 $19.91$ 61.48 51.63
@@ 7 Conclusion
We have introduced the Compliance-to-Code Dataset and a novel structured synthesis pipeline, embodied in our FinCheck system, for executable compliance checking in the complex domain of financial regulation. Our approach, which formalizes regulatory logic as annotated Compliance Units and translates these into verifiable Python code, demonstrates significant gains in functional accuracy (achieving up to 78.1% Pass@1 with gold structures and reasoning steps), interpretability through traceable commentary, and inherent auditability compared to end-to-end LLM prompting. This research not only provides a valuable new resource and strong benchmark results but also opens new avenues for developing safer, more scalable, and reliable RegTech solutions.
@@
Societal Impact: Our approach can streamline compliance and enhance transparency, but should not replace expert judgment in critical situations. We have carefully reviewed and redacted data for release, and recommend that our tools be used with responsible human oversight. For more discussion, please refer to Ethics Statements in Appendix E.
Appendix A Acknowledgements
We would like to express our deepest gratitude to Prof. Zixuan Yuan, our corresponding author, for his exceptional wisdom, visionary guidance, and steadfast mentorship throughout the project. We are grateful to Dr. Yutao Huang (Peking University) for policy consulting, and to Yuting Qiu for legal advice. We thank Siyuan Li for leading the dataset construction and experiments, ensuring data quality and experimental accuracy. Jian Chen served as the project lead and contributed to the overall research framework, dataset and experimental design, manuscript preparation, and initial project coordination. Rui Yao contributed to both dataset construction and experimental implementation. Weihua Qiu, Simin Zhang, Chucheng Dong, and Zhiyao Li assisted in dataset creation. Xuming Hu, Peilin Zhou and Qipeng Xie contributed to the revision of the manuscript. We thank Qipeng Xie for providing computational resources.
This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0630) and NSFC (No. 62402413).
Appendix B License
The Compliance-to-Code dataset is publicly available under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.
@@ Appendix C Accessibility
Appendix D Data Usage
@@
Appendix E Ethics Statements
@@ Appendix H Prompt of Instruction-tuning
Because of the length of the prompt, we only display the structure of the prompt in Figure 4 For more detail, please refer to the dataset.